OpenStudy (anonymous):

Solve for X. 7/sqrt(33) + sqrt(33)/ 4 = 5x

zepdrix (zepdrix):

Hey steph! \c:/ \[\Large \color{royalblue}{\text{Wecome To OpenStudy! :)}}\]

OpenStudy (anonymous):

thanks

zepdrix (zepdrix):

\[\Large \frac{7}{\sqrt{33}}+\frac{\sqrt{33}}{4}=5x\]Mmmmmm, I don't like fractions. Let's multiply both sides by 4,\[\Large \frac{7\cdot4}{\sqrt{33}}+\frac{\sqrt{33}\cdot\cancel4}{\cancel4}=5x\cdot4\]Giving us,\[\Large \frac{28}{\sqrt{33}}+\sqrt{33}=20x\]Now let's multiply both sides by sqrt{33},\[\Large \frac{28\cdot\cancel{\sqrt{33}}}{\cancel{\sqrt{33}}}+\sqrt{33}\cdot\sqrt{33}=20x\cdot\sqrt{33}\]Giving usssssssss,\[\Large 28+33=\left(20\sqrt{33}\right)x\]From here it will be a tad easier to solve for x, now that we've dealt with the fractions. Confused by any of that? :O

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

wouldnt it end up being someting like

OpenStudy (anonymous):

61/22sqrt33=x

zepdrix (zepdrix):

Dividing both sides by the coefficient on x got you there? Yes good! :)\[\Large \frac{61}{20\sqrt{33}}=x\]

zepdrix (zepdrix):

Hmm it's usually improper to leave a root in the denominator like that. Let's multiply the top and bottom by sqrt33.

zepdrix (zepdrix):

\[\Large \frac{61}{20\sqrt{33}}\cdot\frac{\sqrt{33}}{\sqrt{33}}=x\]Which simplifies to,\[\Large \frac{61\sqrt{33}}{20\cdot33}=x\]Understand what I did with that step there? :O

zepdrix (zepdrix):

I mean unless you have multiple choice to choose from. In which case you should just stop when your answer matches :)

OpenStudy (anonymous):

A.\[A.\frac{ 35+5\sqrt{13} }{ 6 }\] B\[\frac{ 7+\sqrt{13} }{ 30 }\] C.\[\frac{ 7+6\sqrt{13} }{ 210 }\] D.\[\frac{ 35+30\sqrt{13} }{ 42 }\] Those are the answer choices.

zepdrix (zepdrix):

Mmm Ok lemme see if I made a boo boo somewhere.

zepdrix (zepdrix):

Are you sure you've matched up the multiple choice to the right question? Because we shouldn't end up with sqrt(13), or addition in the numerator either. :(

OpenStudy (anonymous):

wait ill check.

OpenStudy (anonymous):

nvm i got the answer thanks for the help.

zepdrix (zepdrix):

cool